Use the binomial formula to find the coefficient of the t*x21 term in the expansion of (2t-x)25 Х 5
Answer #1

♡ (27-3325 = 25) (2£325 - (25) (27j24(ex3* +(25) (243% (232 -(23)(2+)2 (a3+ -...-(443) (2+1° (30)25 General term: Toti = (-17